I need help solving for N. I can guess the right answer (which is 4) and use the "plug in and try" method, but I need to know the right steps to solve it. We do not have a math textbook at school. Teacher writes on board and then hands out homework sheets, so I have nothing to refer back to when it comes time to do homework!
Problem:
44 divided by N = 15 minus N

Thank you kindly. My mom is usually very good at math and algebra, but for some reason tonight she just can't work it out!

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

44/N=15-N NOW CROSS MULTIPLY.
N(15-N)=44
15N-N^2=44
-N^2+15N-44=0
N^2-15N+44=0
(N-11)(N-4)=0
N-11=0
N=11 ANSWER.
N-4=0
N=4 ANSWER.

Answer by stanbon(75887)

44 divided by N = 15 minus N
44/N = 15-N
Multiply thru by N to get:
44 = 15N-N^2
Rearrange:
N^2-15N+44 = 0
Factor:
(N-11)(N-4)=0
N=11 or N=4
==================
Cheers,
Stan H.